OpenStudy (anonymous):

find the equation of the plane which passes through P(2,1,-4) and which is perpendicular to both x+y+z=2 and x-y-z=4

OpenStudy (anonymous):

The two equations are equations of planes x + y + z = 2 x - y - z = 4 The normal vectors of these planes are < 1, 1, 1 > and <1 ,-1, -1>. The normal vector to the plane which contains these two vectors is < 1, 1, 1 > X < 1 , -1, -1> = < 0, 1, -1 > (Compute dot products to check this). So the equation of that plane is y - z = const P is in the plane so 1 - -4 = 5 We have 0x + y - z = 5